OpenStudy (anonymous):

Show, without graphing, that the limit as x approaches zero from the positive side of natural log x over x is negative infinity.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0^{+}}\frac{ lnx }{ x }=-\infty \]

OpenStudy (anonymous):

What is posted abouve is what one of my students answered.

OpenStudy (anonymous):

What would you do to answer it?

OpenStudy (anonymous):

1. the domain of the function is x>0 2. denominator is always +ve 3. for numberator, \(\ln x=0\) for \(x=1\). for \(x>1,\quad\ln x>0\) for \(0<x<1,\quad \ln x<0\) 4. thus,in the interval \(0<x<1\), the limit is -ve. then you throuw in a graph